OpenStudy (anonymous):

which expression is equivalent to cot t sec t? A. csc t B. sin t C. tan t D. sin^2 t E. csc^2 t

Nnesha (nnesha):

write sec and cot in terms of sin or cosine :-) sec = ? cot = ?

OpenStudy (anonymous):

what?

Nnesha (nnesha):

what's the definition of sec like tan = sin/cos sin = 1/csc sec = ?

OpenStudy (anonymous):

idk

Nnesha (nnesha):

alright that's the first thing to know for trig \[\tan = \frac{ \sin }{ \cos }~~~~~\cot=\frac{ \cos }{ \sin } ~~~\sin=\frac{ 1 }{ \csc}~~~\cos=\frac{ 1 }{ \sec } \] tan is reciprocal of cos and same with sin and cos so csc = 1/sin sec = ?